# Guarding the Guardrails: A Taxonomy-Driven Approach to Jailbreak Detection

### v1

@@ Abstract
Jailbreaking techniques pose a significant threat to the safety of Large Language Models (LLMs). Existing defenses typically focus on single-turn attacks, lack coverage across languages, and rely on limited taxonomies that either fail to capture the full diversity of attack strategies or emphasize risk categories rather than the jailbreaking techniques. To advance the understanding of the effectiveness of jailbreaking techniques, we conducted a structured red-teaming challenge. The outcome of our experiments are manifold. First, we developed a comprehensive hierarchical taxonomy of 50 jailbreak strategies, consolidating and extending prior classifications into seven broad families, including impersonation, persuasion, privilege escalation, cognitive overload, obfuscation, goal conflict, and data poisoning. Second, we analyzed the data collected from the challenge to examine the prevalence and success rates of different attack types, providing insights into how specific jailbreak strategies exploit model vulnerabilities and induce misalignment. Third, we benchmark a popular LLM for jailbreak detection, evaluating the benefits of taxonomy-guided prompting for improving automatic detection. Finally, we compiled a new Italian dataset of 1364 multi-turn adversarial dialogues, annotated with our taxonomy, enabling the study of interactions where adversarial intent emerges gradually and succeeds in bypassing traditional safeguards.
@@ 1 Introduction
Despite efforts to align LLMs with human preferences through Supervised Fine-Tuning (SFT), often followed by Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2020); Stiennon et al. (2020); Ouyang et al. (2022) or Direct Preference Optimization (DPO) Rafailov et al. (2024), these models can still generate unsafe content, even in response to benign user inputs. As shown by Betley et al. (2025), even small perturbations in fine-tuning, such as a single epoch of training on insecure code, can lead to significant misalignment. These risks are further amplified by adversarial attacks, where malicious actors exploit the model’s vulnerabilities to induce harmful outputs.
@@
To mitigate the risks of misalignment and jailbreaking, guardrailing systems are built as safety layers around the core language model. These systems monitor, constrain, or intervene in the model’s behavior to prevent undesired outputs. Common components include anomaly detectors, prompt sanitizers, decoding constraints, and other filters Jain et al. (2023); Cao et al. (2024); Zeng et al. (2024). Among these, external safety modules play a central role. Examples include content moderation tools, such as the OpenAI Content Moderation API, [^1] https://platform.openai.com/docs/guides/moderation/overview Perspective API, [^2] https://perspectiveapi.com/ and Llama Guard Inan et al. (2023). These detectors are typically implemented as trained classifiers or specialized LLMs fine-tuned on safety-related data to recognize and block malicious activity before harm occurs.
While guardrailing systems provide essential protective layers, their reliability depends heavily on the robustness of their individual components. In particular, adversarial attack detectors must be thoroughly examined before being integrated into safety infrastructures. The more accurate and generalizable they are, the better they can proactively identify emerging jailbreak attempts. To be effective, such systems must cover a broad spectrum of attack strategies across domains and languages. Existing defenses lack robustness against multi-turn jailbreaks Li et al. (2024), as they are assessed only on single-turn adversarial prompts, which represents a threat model that fails to reflect real-world dynamics. Training detectors capable of handling multilingual and multi-turn attacks requires curated datasets with annotated adversarial prompting strategies grounded in a comprehensive taxonomy. However, such data are scarce or unavailable for most languages, including Italian.
Moreover, existing taxonomies primarily emphasize the type of harm produced by an attack rather than the prompting technique that generates it. Others are narrowly focused on a single class of attacks, such as persuasion Zeng et al. (2024), and therefore fail to capture the full diversity of jailbreak strategies. In addition, jailbreak approaches evolve rapidly, and taxonomies that are not updated accordingly quickly lose their relevance. These limitations reduce the usefulness of existing taxonomies for annotation, detector training, and targeted mitigation.
To address these gaps, we present an Italian dataset of 1364 unsafe multi-turn dialogues spanning a wide range of jailbreaking techniques, collected through a structured red-teaming challenge. We propose a hierarchical taxonomy comprising 50 jailbreaks that overcomes the limitations of existing taxonomies and use it to annotate the dataset. Leveraging this resource, we analyze the prevalence of different attack types and evaluate the ability of GPT-5 to detect these adversarial behaviors.
Our overall objective is to contribute to the enhancement of guardrailing systems in which an attack detector plays a central role. We present four main contributions:
@@
    We release a new dataset for evaluating the safety and performance of adversarial prompt detectors in Italian. The dataset covers both single-turn and multi-turn jailbreaks and addresses the critical scarcity of such resources in the field. To the best of our knowledge, this is the first dataset that is simultaneously in Italian, multi-turn, and annotated specifically for jailbreak detection. The dataset and the material will be releases upon acceptance.
@@
    We propose a comprehensive taxonomy of jailbreaking techniques against LLMs. This taxonomy aggregates and extends existing categorizations from the literature and is further refined through empirical observations of real attacks collected during dataset construction, resulting in broader and more detailed coverage than previous approaches.
@@
The remainder of this paper is structured as follows. In Section 2, we review the related work. Section 3 outlined our red teaming challenge for dataset construction (3.1) and the taxonomy design (3.2). The results are presented in Section 4, followed by the findings from our use-case experiments in Section 5. Finally, Section 6 concludes the paper and discusses future research directions.
@@ 2 Related work
Public datasets that explicitly label adversarial techniques in multi-turn dialogue are relatively rare. SafeDialBench Cao et al. (2025) and Multi-Turn Human Jailbreaks (MHJ) Li et al. (2024) offer public multi-turn datasets annotated with 7 distinct jailbreaking techniques each. In MHJ, these labels are informed by red teamers’ own metadata describing their rationale and strategy. Both resources contribute into adversarial prompting in multi-turn setting, yet their restricted taxonomies and focus on English (with Chinese in SafeDialBench) underscore the continued lack of wider coverage.
@@
Narrow-focused studies analyze specific jailbreak strategies in depth. For instance, Wei et al. (2023) identify two key alignment failure modes exploited by jailbreaks: competing objectives and mismatched generalization, while Zeng et al. (2024) examine persuasion mechanisms as a targeted attack vector. In contrast, broader generalization efforts aim to systematize a wider variety of techniques. HackAPrompt Schulhoff et al. (2023) stands out as a key foundation, collecting 600K adversarial prompts through a jailbreaking competition. In a similar direction, Liu et al. (2024) propose a taxonomy of ten techniques organized into three families. Yu et al. (2024b) categorize jailbreak prompts into five categories and ten patterns, grounding their analysis in strategies observed among real users. Finally, Rao et al. (2024) link specific prompting techniques to the underlying adversarial intent. While these taxonomies provide valuable perspectives, none achieve full coverage of the diverse jailbreak strategies observed in practice.
@@ 3.1 Red teaming challenge
Participants were divided into groups and assigned tasks corresponding to one of three vulnerability areas: Attacks on Data, Attacks on the Model, and Attacks on Infrastructure. Across these areas, nine tasks were defined. Attacks on Data included eliciting gender bias, eliciting ethnicity bias, and inducing privacy violations. Attacks on the Model comprised generating information that could cause physical or non-physical harm and triggering hallucinations. Finally, Attacks on Infrastructure involved bypassing arbitrary restrictions defined by the system prompt, revealing a hidden word, and extracting fragments of the prompt itself. To support the infrastructure-related tasks, the system prompt was modified with explicit prohibitions, such as arbitrary restrictions, requiring that the hidden word remains undisclosed, and forbidding any disclosure of the prompt itself.
@@ 3.2 Taxonomy
We present a comprehensive taxonomy of 50 prompt-based jailbreaking techniques targeting Large Language Models, illustrated in Figure 1. This taxonomy consolidates and extends prior classifications from the literature, integrating insights from existing taxonomies and further refining them through empirical observations of attacks collected during our red-teaming challenge.
The taxonomy is organized into three hierarchical levels and groups techniques into seven distinct families, each defined by the primary mechanism through which adversarial prompts bypass safety safeguards: Impersonation Attacks & Fictional Scenarios, Privilege Escalation, Persuasion, Cognitive Overload & Attention Misalignment, Encoding & Obfuscation, Goal-Conflicting Attacks, and Data Poisoning Attacks. In the following subsections, we discuss each family in detail, outlining its underlying mechanisms and representative jailbreak strategies.
@@ 3.2.1 Impersonation Attacks & Fictional Scenarios
This family of attacks induces the model to assume roles or operate within fictional contexts that relax its alignment constraints. This pattern is widely studied in the literature and corresponds to categories found in several taxonomies, including Cognitive Hacking (COG) Rao et al. (2024), Pretending family Liu et al. (2024), and Virtualization category Kang et al. (2023). Instances include:
  • Role Play: The model is prompted to act as specific individuals (e.g., a malicious expert, a criminal, or an unfiltered AI), creating implicit associations between certain roles and unmoderated behavior. Further subdivisions are identified by Yu et al. (2024b), which distinguish between Defined Personas and Virtual AI, with the latter further split into Superior, Opposite, and Alternate modes.
  • Benign Context Framing: Malicious requests are embedded within carefully crafted scenarios that appear to carry lower perceived risk. This includes, for instance, presenting the request as part of academic research, sociological analysis, or a controlled experiment. This category corresponds to the Disguised Intent patterns described in Yu et al. (2024b). A specific variant frames the request within a Game scenario, which in our taxonomy is treated as a third-level subcategory.
  • Fictional Framing: Harmful requests are presented within jokes, stories, or imagined scenarios, making them appear legitimate and creative. This category maps to Imagined Scenario and partially overlaps with Joking Pretext in Yu et al. (2024b).
@@ 3.2.2 Privilege Escalation
Privilege escalation attacks simulate elevated privileges or unconstrained execution contexts to induce the model to bypass its safety restrictions. Typical techniques include claiming administrative roles, declaring that the model has been jailbroken, or using formatting cues that reinforce perceived command authority. This class of attacks corresponds to the one in Liu et al. (2024) and includes:
@@ 3.2.3 Persuasion
Zeng et al. (2024) present a comprehensive analysis of adversarial persuasion, identifying forty techniques grouped into thirteen strategies. Building on this framework and including observations from other works, we distill the strategies most directly relevant to jailbreaks into six primary second-level categories within the Persuasion family.
@@ 3.2.4 Cognitive Overload & Attention Misalignment
These attacks bypass moderation by creating complex or overwhelming contexts that divert the attention of the model away from safety constraints. They exploit both computational and attentional limitations. This class corresponds to the Attention Shifting category described by Yu et al. (2024b).
@@ 3.2.5 Encoding & Obfuscation
When attackers maximize the distance between their requests and the distributions seen during safety training, models may become increasingly vulnerable to unsafe behavior. Wei et al. (2023) describe this phenomenon as Mismatched Generalization. Comparable concepts appear in other taxonomies under different labels, including Orthographic Techniques Rao et al. (2024), Obfuscation Kang et al. (2023), and Character-Level Encoding Liu et al. (2024). Instances include:
@@
  • Linguistic Encoding: Transliteration of the request using alternate representations. This includes low-resource languages, alternative scripts (e.g., Cyrillic look-alikes), emojis, Base64, or other encoding schemes.
@@ 3.2.6 Goal-Conflicting Attacks
  • Prefix Injection: Malicious prefixes are prepended to the prompt so that the model interprets them as part of its conversational history Wei et al. (2023).
@@ 3.2.7 Data Poisoning Attacks
  • Incremental Poisoning: The malicious request is distributed across multiple turn, progressively increasing in harmfulness, often starting with innocent prompts.
@@ 4 Challenge data analysis
Our dataset, collected during the Red Teaming challenge, consists of 1364 unique branches of adversarial conversations, of which 185 correspond to successful attacks. On average, each user-assistant interaction contains 2.85 messages and 176.51 words. For successful cases, these averages slightly change to 3.02 messages and 168.57 words per conversation. This corpus allows us to derive insights of the effectiveness of different adversarial prompting strategies, analyzed here by level.
@@
Figure 2: Distribution of adversarial dialogues across first-level jailbreak families, showing total occurrences and successful attacks for each category.
4.1 First-level jailbreak techniques
@@
Figure 2 illustrates the distribution of adversarial dialogues across the first-level jailbreak categories, showing both the number of occurrences and corresponding successful cases. Detailed results, including success rates, are reported in Table 1.
@@
The most prevalent jailbreak family employed during the red-teaming challenge was Impersonation Attacks & Fictional Scenarios, which appeared in 696 dialogues (51.0% of the total). The Data Poisoning Attacks family achieved the highest success rate (17.2%), while Encoding & Obfuscation techniques showed the lowest (9.4%), having minimal effect on the tested model except for the Lexical Techniques subcategory. Figure 2 and Table 1 also include two auxiliary categories: No Technique and Automated Attacks. The No Technique category accounts for cases in which participants successfully completed jailbreak tasks without applying any explicit attack strategy, directly issuing request to the model. Including this category highlights how the use of targeted jailbreak techniques significantly increases the overall success rate of adversarial attempts.
@@
Finally, the Automated Attacks category represents an orthogonal dimension relative to our taxonomy. In our red teaming challenge, participants could rely only on adversarial prompting and not on optimization-based methods, given the limited time and resources available. However, automatically discovered triggers, previously identified in other studies as transferable across models Zou et al. (2023), were permitted for testing. These Automatic Attacks achieved the highest success rate among all categories, as reported in Table 1.
4.2 Second-level jailbreak techniques

### v2

@@ Abstract
Jailbreaking techniques pose a significant threat to the safety of Large Language Models (LLMs). Existing defenses typically focus on single-turn attacks, lack coverage across languages, and rely on limited taxonomies that either fail to capture the full diversity of attack strategies or emphasize risk categories rather than jailbreaking techniques. To advance the understanding of the effectiveness of jailbreaking techniques, we conducted a structured red-teaming challenge. The outcomes of our experiments are fourfold.
First, we developed a comprehensive hierarchical taxonomy of jailbreak strategies that systematically consolidates techniques previously studied in isolation and harmonizes existing, partially overlapping classifications with explicit cross-references to prior categorizations. The taxonomy organizes jailbreak strategies into seven mechanism-oriented families: impersonation, persuasion, privilege escalation, cognitive overload, obfuscation, goal conflict, and data poisoning. Second, we analyzed the data collected from the challenge to examine the prevalence and success rates of different attack types, providing insights into how specific jailbreak strategies exploit model vulnerabilities and induce misalignment. Third, we benchmarked GPT-5 as a judge for jailbreak detection, evaluating the benefits of taxonomy-guided prompting for improving automatic detection. Finally, we compiled a new Italian dataset of 1364 multi-turn adversarial dialogues, annotated with our taxonomy, enabling the study of interactions where adversarial intent emerges gradually and succeeds in bypassing traditional safeguards.
@@ 1 Introduction
Despite efforts to align LLMs with human preferences through Supervised Fine-Tuning (SFT), often followed by Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2020); Stiennon et al. (2020); Ouyang et al. (2022) or Direct Preference Optimization (DPO) Rafailov et al. (2024), these models can still generate unsafe content, even in response to benign user inputs. As shown by Betley and others (2025), even small perturbations in fine-tuning, such as a single epoch of training on insecure code, can lead to significant misalignment. These risks are further amplified by adversarial attacks, where malicious actors exploit the model’s vulnerabilities to induce harmful outputs.
@@
To mitigate the risks of misalignment and jailbreaking, guardrailing systems are built as safety layers around the core language model. These systems monitor, constrain, or intervene in the model’s behavior to prevent undesired outputs. Common components include anomaly detectors, prompt sanitizers, decoding constraints, and other filters Jain and others (2023); Cao and others (2024); Zeng and others (2024). Among these, external safety modules play a central role. Examples include content moderation tools, such as the OpenAI Content Moderation API, [^1] https://platform.openai.com/docs/guides/moderation/overview Perspective API, [^2] https://perspectiveapi.com/ and Llama Guard Inan and others (2023). These detectors are typically implemented as trained classifiers or specialized LLMs fine-tuned on safety-related data to recognize and block malicious activity before harm occurs.
While guardrailing systems provide essential protective layers, their effectiveness depends on the accuracy and generalizability of adversarial attack detectors. To be effective in practice, such systems must cover a broad spectrum of attack strategies across domains and languages. This limitation is reinforced by the fact that many deployed guardrails rely on risk-based taxonomies, that concentrate on what unsafe content is produced rather than capturing how adversarial prompting steers the model. In multi-turn attacks, malicious intent is often distributed across benign-looking turns and only becomes explicit late in the interaction because of the accumulation of harmful context, so harm-based labels provide limited signals for detecting the intermediate steering steps that make these jailbreaks succeed.
Existing defenses lack robustness against multi-turn jailbreaks Li et al. (2024), as they are assessed only on single-turn adversarial prompts, which represents a threat model that fails to reflect real-world dynamics. Training detectors capable of handling multilingual and multi-turn attacks requires curated datasets with annotated adversarial prompting strategies grounded in a comprehensive taxonomy. However, such data are scarce or unavailable for most languages, including Italian. In this work we present four main contributions:
@@
    We release a new dataset for evaluating the safety and performance of adversarial prompt detectors in Italian. The dataset covers both single-turn and multi-turn jailbreaks and addresses the critical scarcity of such resources in the field. To the best of our knowledge, this is the first dataset that is simultaneously annotated specifically for jailbreak detection and multi-turn.
@@
    We propose a multi-level and mechanism-oriented taxonomy of jailbreaking techniques against LLMs. Rather than introducing novel attack methods, our contribution lies in systematically interrelating, consolidating, and harmonizing techniques that prior works have examined in isolation. The taxonomy aligns and integrates these techniques through standardized terminology and explicit cross-references to existing categorizations, while refining their formalization based on recurrent adversarial patterns empirically observed during data collection. To the best of our knowledge, no existing work provides a taxonomy that is simultaneously comprehensive in coverage, mechanism-oriented, multi-level, and empirically validated.
@@
All the materials, including the dataset, are made publicly available at https://lab-rococo-sapienza.github.io/Guarding/.
The remainder of this paper is structured as follows. In Section 2, we review the related work. Section 3 outlines our red teaming challenge for dataset construction (3.1) and the taxonomy design (3.2). The results are presented in Section 4, followed by the findings from our use-case experiments in Section 5. Finally, Section 6 concludes the paper and discusses future research directions.
@@ 2 Related work
Public datasets that explicitly label adversarial techniques in multi-turn dialogue are relatively rare. SafeDialBench Cao et al. (2025) and Multi-Turn Human Jailbreaks (MHJ) Li et al. (2024) offer public multi-turn datasets annotated with 7 distinct jailbreaking techniques each. In MHJ, these labels are informed by red teamers’ own metadata describing their rationale and strategy.
@@
Narrow-focused studies analyze specific jailbreak strategies in depth. For instance, Wei et al. (2023) identify two key alignment failure modes exploited by jailbreaks: competing objectives and mismatched generalization, while Zeng and others (2024) examine persuasion mechanisms as a targeted attack vector. In contrast, broader generalization efforts aim to systematize a wider variety of techniques. HackAPrompt Schulhoff et al. (2023) stands out as a key foundation, collecting 600K adversarial prompts through a jailbreaking competition. In a similar direction, Liu et al. (2024b) propose a taxonomy of ten techniques organized into three families. Yu et al. (2024c) categorize jailbreak prompts into five categories and ten patterns, grounding their analysis in strategies observed among real users. Finally, Rao et al. (2024) link specific prompting techniques to the underlying adversarial intent. While these taxonomies provide valuable perspectives, none achieve full coverage of the diverse jailbreak strategies observed in practice.
@@ 3.1 Red teaming challenge
Participants were divided into groups and assigned tasks corresponding to one of three vulnerability areas: Attacks on Data, Attacks on the Model and Attacks on Infrastructure. Across these areas, nine tasks were defined. Attacks on Data included eliciting gender bias, eliciting ethnicity bias, and inducing privacy violations. Attacks on the Model comprised generating information that could cause physical or non-physical harm and triggering hallucinations. Finally, Attacks on Infrastructure involved bypassing arbitrary restrictions defined by the system prompt, revealing a hidden word, and extracting fragments of the prompt itself. To support the infrastructure-related tasks, the system prompt was modified with explicit prohibitions, such as arbitrary restrictions, requiring that the hidden word remains undisclosed, and forbidding any disclosure of the prompt itself.
@@ 3.2 Taxonomy
We present a comprehensive taxonomy of prompt-based jailbreaking techniques targeting Large Language Models, illustrated in Figure 1. This taxonomy consolidates and extends prior classifications from the literature, integrating insights from existing taxonomies and further refining them through empirical observations of attacks collected during our red-teaming challenge.
The taxonomy is organized into three hierarchical levels and groups techniques into seven distinct families, each defined by the primary mechanism through which adversarial prompts bypass safety safeguards: Impersonation Attacks & Fictional Scenarios, Privilege Escalation, Persuasion, Cognitive Overload & Attention Misalignment, Encoding & Obfuscation, Goal-Conflicting Attacks, and Data Poisoning Attacks. In the following subsections, we discuss each family in detail, outlining its underlying mechanisms and representative jailbreak strategies. See Appendix A for concrete examples of each jailbreaking technique.
@@ 3.2.1 Impersonation Attacks & Fictional Scenarios
This family of attacks induces the model to assume roles or operate within fictional contexts that relax its alignment constraints. This pattern is widely studied in the literature and corresponds to categories found in several taxonomies, including Cognitive Hacking (COG) Rao et al. (2024), Pretending family Liu et al. (2024b), and Virtualization category Kang et al. (2023). Instances include:
  • Role Play: The model is prompted to act as specific individuals (e.g., a malicious expert, a criminal, or an unfiltered AI), creating implicit associations between certain roles and unmoderated behavior (see Character Role Play in Liu et al. (2024b)). Further subdivisions are identified by Yu et al. (2024c), which distinguish between Defined Personas and Virtual AI, with the latter further split into Superior, Opposite, and Alternate modes.
  • Benign Context Framing: Malicious requests are embedded within carefully crafted scenarios that appear to carry lower perceived risk. This includes, for instance, presenting the request as part of academic research, sociological analysis, or a controlled experiment. This category corresponds to the Disguised Intent patterns described in Yu et al. (2024c). A specific variant frames the request within a Game scenario, which in our taxonomy is treated as a third-level subcategory.
  • Fictional Framing: Harmful requests are presented within jokes, stories, or imagined scenarios, making them appear legitimate and creative. This category maps to Imagined Scenario and partially overlaps with Joking Pretext in Yu et al. (2024c).
@@ 3.2.2 Privilege Escalation
Privilege escalation attacks simulate elevated privileges or unconstrained execution contexts to induce the model to bypass its safety restrictions. Typical techniques include claiming administrative roles, declaring that the model has been jailbroken, or using formatting cues that reinforce perceived command authority. This class of attacks corresponds to the one in Liu et al. (2024b) and includes:
@@ 3.2.3 Persuasion
Zeng and others (2024) present a comprehensive analysis of adversarial persuasion, identifying forty techniques grouped into thirteen strategies. Building on this framework and including observations from other works, we distill the strategies most directly relevant to jailbreaks into six primary second-level categories within the Persuasion family.
@@ 3.2.4 Cognitive Overload & Attention Misalignment
These attacks bypass moderation by creating complex or overwhelming contexts that divert the attention of the model away from safety constraints. They exploit both computational and attentional limitations. This class corresponds to the Attention Shifting category described by Yu et al. (2024c).
@@ 3.2.5 Encoding & Obfuscation
When attackers maximize the distance between their requests and the distributions seen during safety training, models may become increasingly vulnerable to unsafe behavior. Wei et al. (2023) describe this phenomenon as Mismatched Generalization. Comparable concepts appear in other taxonomies under different labels, including Orthographic Techniques Rao et al. (2024), Obfuscation Kang et al. (2023), and Character-Level Encoding Liu et al. (2024b). Instances include:
@@
  • Linguistic Encoding: Transliteration of the request using alternate representations. This includes low-resource languages, alternative scripts (e.g., Cyrillic look-alikes), emojis, Base64, or other encoding schemes. The use of low-resource languages can significantly reduce the effectiveness of safety filters Al Ghanim et al. (2024); Deng et al. (2024). In our taxonomy, this case is treated separately from its parent category.
@@ 3.2.6 Goal-Conflicting Attacks
  • Prefix Injection: Malicious prefixes are prepended to the prompt so that the model interprets them as part of its conversational history Wei et al. (2023). A common variant requires the model to begin its answer with a specific phrase. Such attacks exploit the model’s tendency to preserve conversational coherence.
@@ 3.2.7 Data Poisoning Attacks
  • Incremental Poisoning: The malicious request is distributed across multiple turn, progressively introducing problematic elements increasing in harmfulness, often starting with innocent prompts.
@@ 4 Challenge data analysis
Our dataset, collected during the Red Teaming challenge, consists of 1364 unique branches of adversarial conversations, of which 185 correspond to successful attacks. On average, each user-assistant interaction contains 2.85 messages and 176.51 words. For successful cases, these averages slightly change to 3.02 messages and 168.57 words per conversation.
@@
Figure 2: Distribution of adversarial dialogues across jailbreak families, showing total occurrences and successful attacks for each category.
4.1 Distribution of jailbreak families
@@
Figure 2 illustrates the distribution of adversarial dialogues across the seven first-level jailbreak families in our taxonomy, reporting both total occurrences and successful cases. Detailed results, including success rates, are reported in Table 1.
@@
The most prevalent jailbreak family employed during the red-teaming challenge was Impersonation Attacks & Fictional Scenarios, which appeared in 696 dialogues (51.0% of the total). The Data Poisoning Attacks family achieved the highest success rate (17.2%), while Encoding & Obfuscation techniques showed the lowest (9.4%).
Figure 2 and Table 1 also include two auxiliary categories: No Technique and Automatic Attacks. The No Technique category accounts for cases in which participants successfully completed jailbreak tasks without applying any explicit attack strategy, directly issuing requests to the model. Including this category highlights how the use of targeted jailbreak techniques significantly increases the overall success rate of adversarial attempts.
@@
Finally, the Automatic Attacks category represents an orthogonal dimension relative to our taxonomy. In our red teaming challenge, participants could rely only on adversarial prompting and not on optimization-based methods, given the limited time and resources available. However, automatically discovered triggers, previously identified in other studies as transferable across models Zou et al. (2023), were permitted for testing. When an attack relied primarily on such pre-optimized triggers, we annotated it under Automatic Attacks rather than attributing it to one of the seven mechanism-oriented families. As shown in Table 1, this auxiliary category achieves the highest success rate overall.
4.2 Analysis of jailbreak techniques